If the distance between earth and he ⛅ sun is increased by 2%, then find the  percentage change on the gravitational force acting between them???

Grade:12th pass

1 Answers

Arun
6 years ago
Distance between sun and earth= r
mass of earth= m
mass of sun= M
Force of attraction = F= GmM/r²
r' = r + 0.02r
r'= 102/100r

F'= GmM/r'²

F'/F= GmM/r'² / GmM/r² = (r/r' )²
= 100/102)²
1-F'/F = 1- (100/102)²
decrease in F= 1- 0.96
= 0.039 or 3.96%